# Channel Balance Interpolation in the Lightning Network via Machine Learning thanks: * Jointly supervised

### v1

@@ Abstract
The Bitcoin Lightning Network is a Layer 2 payment protocol that addresses Bitcoin’s scalability by facilitating quick and cost-effective transactions through payment channels. This research explores the feasibility of using machine learning models to interpolate channel balances within the network, which can be used for optimizing the network’s pathfinding algorithms. While there has been much exploration in balance probing and multipath payment protocols, predicting channel balances using solely node and channel features remains an uncharted area. This paper evaluates the performance of several machine learning models against two heuristic baselines and investigates the predictive capabilities of various features. Our model performs favorably in experimental evaluation, outperforming by 10% against an equal split baseline where both edges are assigned half of the channel capacity.
Index Terms:
Bitcoin, Lightning Network, Pathfinding, Machine Learning, Payment Channel Network
@@ I Introduction
The Lightning Network is a layer-2 protocol on the Bitcoin blockchain that enables rapid and cost-effective payments. Central to optimizing its efficiency is the accurate estimation of channel balances, which directly informs pathfinding strategies. Effective pathfinding circumvents the need for a cumbersome trial-and-error process in identifying viable payment routes, thereby streamlining operations within the network. While there have been advances in understanding balance generation[1], along with the exploration of reinforcement learning (RL) for pathfinding[2] and the development of multipath payment protocols[3], the prediction of channel balances is less explored territory.
This gap in the literature motivates the current study, which seeks to determine whether it is feasible to interpolate the balances of channels within the Lightning Network accurately. It further aims to identify the most predictive features for balance interpolation, questioning whether these predictions can rely solely on node features, channel features, or a combination of both, and possibly enhanced by topological information of the network. To tackle these questions, the study evaluates the performance of two baseline methods against six machine learning (ML) models, each with a varying subset of features.
The investigation reveals that ML models are indeed capable of predicting channel balances with a better degree of accuracy than heuristic methods like the equal split assumption. This outcome suggests that the predictive power of machine learning models could be leveraged to refine pathfinding algorithms. This study sheds light on the potential of specific features and how machine learning can enhance the network’s pathfinding efficiency.
The remainder of this paper is organized as follows. Section II provides background information on Bitcoin and the Lightning Network. Section III elaborates on the specific problem and the motivation for improved channel balance interpolation. Section IV reviews related work in areas such as pathfinding algorithms and multipath payment protocols for the Lightning Network. Sections V,VI, and VII formally define the problem statement, the data collection and preprocessing steps, and the machine learning models and features evaluated. Section VIII outlines the methodology and metrics. Sections IX and X present and analyze the performance of the various models. Finally, Section XI explores potential future work, such as implementing an enhanced pathfinding algorithm that leverages the balance interpolation model and directions for further improving model performance.
@@ II-A Bitcoin
[…]
@@ II-B Lightning Network
In contrast, the current balance of the channel is contained in the commitment transaction. Both peers maintain a copy of this transaction, but defer its broadcast until they want to close the channel. When payments are made via this channel, the peers jointly update the commitment transaction. Because the peers defer broadcasting the commitment transaction, payments on the Lightning Network are not subject to the 10 minute average blocktime on Layer 1 (the Bitcoin blockchain). Furthermore, the pricing dynamic of LN is fundamentally different than on L1. While transactions on L1 compete for blockspace by paying higher fees to miners, routing nodes on LN compete with each other to relay payments by offering lower fees.
Of course there are tradeoffs to using LN over L1. Single path payments can only be relayed if there exists a path with sufficient liquidity from source to destination. In other words, one cannot forward a payment if they have an insufficient balance on their side of the channel. However, only the nodes incident to a channel know the balance of that channel. Thus payment pathfinding in the Lightning Network is a trial-and-error process.
@@ III Motivation
Topological analysis from other works have shown that the Lightning Network topology is consistent with small-world architecture[7]. These graphs exhibit a robustness because of their high local clustering coefficient and low average path length. However, in the Lightning Network, this abundance of short paths comes with the downside of not knowing which ones have sufficient liquidity to relay the payment. For example, River is a Bitcoin-only exchange that operates two of the largest nodes in the Lightning Network[8, 9]. In a research report River released in 2023, they cited timeout during pathfinding as the leading cause of payment failure for their nodes[10]. The goal of this work is to investigate ways of evaluating paths before trying them. Our key focus is on whether the balance of intermediate channels can be interpolated from channel policies and node metrics. This would enable the prioritization of paths in the trial-and-error process, and thus save time searching for a viable path.
@@ IV Related Work
LightningNetworkDaemon, the majority implementation present on the Lightning Network [11] has two models for estimating the success probabilities of paths: apriori and bimodal. The apriori model assumes a constant probability on untested channels also taking into account capacity of the channel and payment size[12]. The bimodal model is an outcome of Pickhardt et al.’s work[1] whereby they sampled balances in the Lightning Network and found that portions of channels followed different distributions. They observed that about 30% of channels balances were concentrated on one side while the rest of the channel balances made up a uniform distribution. The assumption of a bimodal distribution is a configurable path estimator option in LND as of version v0.16.
The next majority implementation present on the Lightning Network is core-lightning. In the default pathfinding behavior, paths are weighted using a fee and risk factor. This weight takes into consideration the risk of locking up liquidity for some number of blocks, with bias to larger capacity channels. The core-lightning implementation also includes a plugin architecture to extend and customize its functionality. One such plugin, altpay, implements pickhardt payments with probability scoring. This payment algorithm is the result of work by Pickhardt and Richter to generalize payments as a minimum cost flow problem[3]. Here, the cost considers both the fee and reliability of a path. The reliability of the path is estimated using an uncertainty network. The uncertainty network is updated in a similar fashion to the trial-and-error approach in a single path payment. However, multipath payments explore multiple paths, are split into smaller amounts, and are not constrained by the liquidity of a single route. In their work, these authors observed that multipath payments can more reliably move funds across the network as compared to single path payments.
Other works include Valko and Kudenko’s work on using reinforcement learning for path planning in the Lightning Network[2]. In this technique, an agent learns to predict paths in the Lightning Network that have low total cost and are close to the shortest path calculated by Dijkstra’s algorithm. This approach has a nearly constant execution time regardless of the number of nodes considered, however, it also requires limited retraining ($<$100 epochs) when the network updates.
@@ V Problem Formulation
Graph $G$ has the constraint that if an edge exists in one direction, it must also exist in the opposite direction, i.e., $(u,v)\in E\iff(v,u)\in E$. The set of two edges between any two nodes is called a channel, denoted as $\{(u,v),(v,u)\}$. For simplicity, we represent a channel by the set of its two nodes: $\{u,v\}$. The total capacity of the channel $\{u,v\}$ is defined as $c_{\{u,v\}}=y_{(u,v)}+y_{(v,u)}$.
@@
Moreover, since we are given $c_{\{u,v\}}$ for all edges, and we know $0\leq y_{(u,v)}\leq c_{\{u,v\}}$, we also have that $y_{(u,v)}=p_{(u,v)}c_{\{u,v\}}$, where $0\leq p_{(u,v)}\leq 1$. Intuitively, $p_{(u,v)}$ is the proportion of the channel capacity which belongs to the $(u,v)$ direction. From this, we see that we focus on predicting $p_{(u,v)}$, and then use it to obtain $y_{(u,v)}$.
@@ VI Data Collection and Preprocessing
The data used in this experiment is a combination of publicly available information from the Lightning Network and crowdsourced information from nodes in the network. A snapshot of the network from December 15, 2023 is used in this experiment. The snapshot was collected from a lightning node on the network operated by Amboss Technologies. Further information is derived from the network snapshot. This information includes aggregated channel capacities, giving the capacity of each node and the network. Other relevant information includes the fee ratios of each node which is found by dividing average incoming fee rates by average outgoing fee rates. In other works[13], this ratio is referred to as drain. Balance data is crowdsourced from nodes in the Lightning Network. Node operators may opt to share balance with Amboss Technologies in order to receive liquidity notifications, automatically purchase channels, or to maintain a historical account.
@@ VI-A Preprocessing Steps
Balance information for each node is represented by its local balance, recorded at one-minute intervals over the preceding hour. This data is converted into a probability density function (PDF) through kernel density estimation. Subsequently, a balance value is sampled from each PDF, serving as the representative local balance for the corresponding channel within the dataset.
@@ VI-B Preliminary Analysis
The following section is a statistical overview of the Lightning Network, including various tables and metrics that detail the distribution of channel capacities and the ratio of channel balances to their capacities. The analysis of channel capacities illustrate the range and distribution of liquidity across the network. This data serves as a reference for understanding the current state and structural characteristics of the Lightning Network.
@@
The leftmost histogram in Figure 3 displays the frequency of channel capacities in the Bitcoin network. Capacities range between 0 and 10 BTC, with the highest frequency observed at the lowest capacity interval. The frequencies diminish significantly for higher capacities, although there are several notable peaks, suggesting capacity values around round numbers are more common. The overall distribution suggests that smaller channel capacities are much more typical within this network, with occasional larger-capacity channels appearing less frequently.
The middle histogram in Figure 3 depicts the distribution of edge balances in Bitcoin (BTC). This data also shows a left-skewed distribution, indicating a prevalence of smaller balances over larger ones.
The rightmost histogram in Figure 3 illustrates the distribution of edge balances as a percentage of their respective capacities within a network. The distribution shows two prominent peaks indicating a high frequency of channels with balances that are either very low (close to empty) or very high (close to full capacity) relative to their channel capacities. There is a notable scarcity of channels with intermediate balances. This pattern could be indicative of a ”bimodal” distribution, suggesting that channels within this network are typically either mostly unused or nearly fully utilized, with fewer channels maintaining a balance around the mid-range. This distribution aligns with findings reported in the study by Pickhardt et al [1].
@@ VI-C Feature Analysis
To assess the relationship between each feature and the target variable, channel balance, a correlation analysis was conducted. Presented below in Figure 4 is a table showcasing those features which demonstrated any notable correlation, in this case, having a correlation coefficient beyond the range of $-0.1$ to $+0.1$. The feature Max HTLC msat represents the maximum size of an outgoing payment a node will handle. Feature flag #19 acts as a binary flag that signifies whether the node permits the creation of channels with capacities exceeding approximately 16.7 million satoshis, which is generally referred to as the ’wumbo channel’ threshold in the Lightning Network.
@@
local_feat_19.is_known 0.1523
remote_feat_19.is_known 0.1603
local_max_htlc_msat 0.6246
remote_max_htlc_msat 0.6165
Figure 4: Table of features and their correlations to the target variable.
Both Max HTLC msat and Feature 19 have a positive correlation with the target variable, as indicated by the positive values of the correlation coefficients. The features local_max_htlc_msat and remote_max_htlc_msat, show a notably strong positive correlation. In contrast, the features local_19.is_known and remote_19.is_known exhibit a modest but still significant positive correlation.
@@ VII-A Modeling
where $\Theta$ are learnable weights, $\mathbf{x}_{u}$, $\mathbf{x}_{v}$ and $\mathbf{e}_{(u,v)}$ are the node and edge features respectively, while $c_{\{u,v\}}$ is the capacity of the channel. While several choices are possible for $f_{\Theta}$, such as multi-layer perceptrons or Graph Neural Networks, we focus on Random Forests for this work given their simplicity and efficacy. In particular, our Random Forest (RF) model operates on the concatenation of the features of the source and destination nodes as well as the edge features:
$\hat{p}_{(u,v)}=\mathrm{RF}(x_{u}||z_{u}||x_{v}||z_{v}||e_{(u,v)})$
The model is trained using a Mean Squared Error loss. Moreover, we know that $p_{(u,v)}=1-p_{(v,u)}$. Therefore, we would like to design our model such that $f_{\Theta}(u,v,G,x,c)=1-f_{\Theta}(v,u,G,x,c)$. We can obtain this approximately through data augmentation: make sure that both directions of an edge are present in the data so that the model is trained to predict both that $f_{\Theta}(u,v,G,x,c)=p_{(u,v)}$ and that $f_{\Theta}(v,u,G,x,c)=1-p_{(v,u)}$.
@@ VII-B Node Features
    The node’s capacity divided by the network’s capacity. This indicates how much of the network’s capacity is incident to a node.
@@
    Ratio of the mean cost of a nodes outgoing fees to the mean cost of its incoming fees.
@@ VII-C Edge Features
  • Min HTLC The minimum amount this edge will route. (denominated in millisats)
  • Max HTLC msat The maximum amount this edge will route. (denominated in millisats)
  • Fee Rate millimsat Proportional fee to route a payment along an edge. (denominated in millimillisats)
  • Fee Base msat Fixed fee to route a payment along an edge. (denominated in millisats)
@@ VII-D Positional Encodings
Positional encoding is essential for capturing the structural context of nodes, as graphs lack inherent sequential order. Utilizing eigenvectors of the graph Laplacian matrix as positional encodings provides a robust solution to this challenge. These eigenvectors highlight key structural patterns, enriching node features with information about the overall topology of the graph[14]. By integrating these spectral properties, machine learning models can effectively recognize and utilize global characteristics, enhancing performance in tasks like node classification and community detection.

### v2

@@ Abstract
The Bitcoin Lightning Network is a Layer 2 payment protocol that addresses Bitcoin’s scalability by facilitating quick and cost-effective transactions through payment channels. This research explores the use of machine learning models to interpolate channel balances within the network, which can be used for optimizing the network’s pathfinding algorithms. While there has been much exploration in balance probing and multipath payment protocols, predicting channel balances using solely node and channel features remains an uncharted area. This paper evaluates the performance of several machine learning models against two heuristic baselines and investigates the predictive capabilities of various features. Our model performs favorably in experimental evaluation, reducing the relative error by 27% compared to an equal split baseline where both edges are assigned half of the channel capacity.
@@ I Introduction
The Lightning Network is a layer-2 protocol on the Bitcoin blockchain that enables rapid and cost-effective payments. Central to optimizing its efficiency is the accurate estimation of channel balances, which directly informs pathfinding strategies. Effective pathfinding circumvents the need for a cumbersome trial-and-error process in identifying viable payment routes. While there have been advances in understanding pathfinding with uncertain channel balances [1], exploration of reinforcement learning (RL) for pathfinding [2] and the development of multipath payment protocols [3], the direct prediction of channel balances is less explored territory.
This gap in the literature motivates the current study, which seeks to determine whether it is feasible to interpolate the balances of channels within the Lightning Network accurately. Furthermore, this work aims to identify the most predictive features for balance interpolation, questioning whether these predictions can rely solely on node features, channel features, or a combination of both, and possibly enhanced by topological information of the network. To tackle these questions, the study evaluates the performance of two baseline methods against six machine learning (ML) models, each with a varying subset of features.
The investigation reveals that ML models can predict channel balances more accurately than heuristic methods like the equal split assumption. This predictive power can be leveraged to refine pathfinding algorithms by reducing uncertainty in channel balances and decreasing failed payment attempts.
@@ II-A Bitcoin
Bitcoin is a peer-to-peer electronic cash system invented in 2008 by a person or group under the pseudonym Satoshi Nakamoto [4]. The purpose of Bitcoin is to provide a trustless way to transfer value. The features which allow this trustless transfer of value are Bitcoin’s distributed ledger, Proof-of-Work consensus model, and fixed supply. The Bitcoin network operates without relying on a central server. New transactions are added to the decentralized ledger through a process known as mining. Bitcoin nodes which participate in this mining process collect transactions and add them to to an “ongoing chain of hash-based proof-of-work” [4], by solving a computationally difficult problem. The chain which demonstrates the most computational effort is regarded as the authentic record of transactions.
@@ II-B Lightning Network
In contrast, the latest balance of the channel is contained in the commitment transaction. Both peers maintain a copy of this transaction but defer its broadcast until they want to close the channel. When payments are made through this channel, the peers jointly update the commitment transaction. Because the peers defer broadcasting the commitment transaction, payments on the Lightning Network are not subject to the 10-minute average blocktime on Layer 1 (the Bitcoin blockchain). Furthermore, the pricing dynamic of LN is fundamentally different from that of L1. While transactions on L1 compete for blockspace by paying higher fees to miners, routing nodes on LN compete with each other to relay payments by offering lower fees.
Of course, there are trade-offs to using LN over L1. Single-path payments can only be relayed if there exists a path with sufficient liquidity from source to destination. In other words, one cannot forward a payment if they have an insufficient balance on their side of the channel. However, only the nodes incident to a channel know the balance of that channel. An example of these concepts can be seen in Figure 2 and Figure 2. Thus, payment pathfinding in the Lightning Network is a trial-and-error process.
@@ III Motivation
Topological analysis from other works has shown that the Lightning Network topology is consistent with the small-world architecture [7]. These graphs exhibit robustness because of their high local clustering coefficient and low average path length. However, in the Lightning Network, this abundance of short paths comes with the downside of not knowing which ones have sufficient liquidity to relay the payment. For example, River is a Bitcoin-only exchange that operates two of the largest nodes in the Lightning Network [8, 9]. In a research report by River published in 2023, they cited timeout during pathfinding as the leading cause of payment failure for their nodes [10]. The purpose of this work is to investigate ways of evaluating paths before trying them. Our key focus is on whether the balance of intermediate channels can be interpolated from channel policies and node metrics. This would enable the prioritization of paths in the trial-and-error process and save time in the search for a viable path.
@@ IV Related Work
LightningNetworkDaemon, the majority implementation present on the Lightning Network [11], has two models for estimating the success probabilities of paths: apriori and bimodal. The apriori model assumes a constant probability on untested channels, taking into account channel capacity and payment size [12]. The bimodal model is an outcome of Pickhardt et al.’s work [1] where they sampled balances in the Lightning Network and found that portions of channels followed different distributions. They observed that about 30% of the channel balances were concentrated on one side while the rest of the channel balances made up a uniform distribution. The assumption of a bimodal distribution is a configurable path estimator option in LND as of version v0.16.
The next majority implementation present on the Lightning Network is core-lightning. In the default pathfinding behavior, paths are weighted using a fee and risk factor. This weight takes into account the risk of locking up liquidity for some number of blocks, with bias to larger capacity channels. The core-lightning implementation also includes a plugin architecture to extend and customize its functionality. One such plugin, altpay, implements pickhardt payments with probability scoring. This payment algorithm is the result of work by Pickhardt and Richter to generalize payments as a minimum-cost flow problem [3]. Here, the cost considers both the fee and the reliability of a path. The reliability of the path is estimated using an uncertainty network. The uncertainty network is updated in a similar fashion to the trial-and-error approach in a single-path payment. However, multipath payments explore multiple paths, are split into smaller amounts, and are not constrained by the liquidity of a single route. The authors found that multipath payments move funds across the network more reliably than single-path payments.
The work of Tikhomirov et. al demonstrates how one can systematically probe channels by leveraging payment errors and response times, effectively revealing balances [13]. However, these methods impose network overhead, require capital, and compromise privacy. The reliance on active interaction makes probing costly and intrusive, as it requires continuous test transactions to infer balances. The capital requirement is two-fold: not only does probing demand upfront liquidity, but it also limits the recoverable balances to amounts smaller than the deployed capital.
Other works include Valko and Kudenko’s work on using reinforcement learning for path planning in the Lightning Network [2]. In this technique, an agent learns to predict paths in the Lightning Network that have low total cost and are close to the shortest path calculated by Dijkstra’s algorithm. However, the paper does not explicitly consider real-world channel balances in the Lightning Network.
@@ V Problem Formulation
Graph $G$ has the constraint that if an edge exists in one direction, it must also exist in the opposite direction, i.e., $(u,v)\in E\iff(v,u)\in E$. The set of two edges between a pair of nodes is called a channel, denoted as $\{(u,v),(v,u)\}$. For simplicity, we represent a channel by the set of its two nodes: $\{u,v\}$. The total capacity of the channel $\{u,v\}$ is defined as $c_{\{u,v\}}=y_{(u,v)}+y_{(v,u)}$.
@@
Moreover, since we are given $c_{\{u,v\}}$ for all edges and we know $0\leq y_{(u,v)}\leq c_{\{u,v\}}$, we also know that $y_{(u,v)}=p_{(u,v)}c_{\{u,v\}}$, where $0\leq p_{(u,v)}\leq 1$. Intuitively, $p_{(u,v)}$ is the proportion of the channel capacity that belongs to the $(u,v)$ direction. From this, we focus on predicting $p_{(u,v)}$, and then use it to obtain $y_{(u,v)}$.
@@ VI Data Collection and Preprocessing
The data used in this experiment is a combination of publicly available information from the Lightning Network and crowdsourced information from nodes in the network. This experiment uses a network snapshot from December 15, 2023. The snapshot was collected from a lightning node on the network operated by Amboss Technologies. Further information is derived from the network snapshot. This information includes aggregated channel capacities, giving the capacity of each node and the network. Other relevant information includes the fee ratios of each node which is found by dividing average incoming fee rates by average outgoing fee rates. In other works [14], this ratio is referred to as drain. Balance data is crowdsourced from nodes in the Lightning Network. Node operators may opt to share balance information with Amboss Technologies in order to receive liquidity notifications, automatically purchase channels, or maintain a historical account.
@@ VI-A Preprocessing Steps
Each node’s balance is represented by its local balance, recorded at one-minute intervals over the past hour. This data is converted into a probability density function (PDF) through kernel density estimation. Subsequently, a balance value is sampled from each PDF, serving as the representative local balance for the corresponding channel within the dataset.
@@ VI-B Preliminary Analysis
The following section is a statistical overview of the Lightning Network, including various tables and metrics that detail the distribution of channel capacities and the ratio of channel balances to their capacities. The analysis of channel capacities highlights the range and distribution of liquidity across the network. This data serves as a reference for understanding the current state and structural characteristics of the Lightning Network.
@@
The leftmost histogram in Figure 3 shows the distribution of channel capacities in the Bitcoin network. Most channels have capacities between 0 and 10 BTC, with the highest concentration at the lowest capacity range. Frequencies decline sharply for larger capacities, though distinct peaks suggest a preference for round-number values. Overall, smaller channel capacities dominate, while larger ones are relatively rare.
The middle histogram in Figure 3 depicts the distribution of known edge balances denominated in Bitcoin (BTC). This data also shows a left-skewed distribution, indicating a prevalence of smaller balances over larger ones.
The rightmost histogram in Figure 3 depicts the distribution of known edge balances as a percentage of their respective capacities. It reveals a bimodal pattern, with peaks at balances near empty and near full capacity, while intermediate balances are notably scarce. This suggests that channels are typically either minimally used or nearly fully utilized, rather than maintaining mid-range balances. These findings align with those reported by Pickhardt et al. [1].
@@ VI-C Feature Analysis
To evaluate the relationship between each feature and channel balance, a correlation analysis was performed. Table I presents features with notable correlations, defined as having a correlation coefficient outside the range of [0.1,+0.1][-0.1,+0.1]. The Max HTLC feature represents the maximum outgoing payment size (denominated in millisatoshis) a node will process. Feature flag #19 is a binary indicator of whether a node allows channels exceeding approximately 16.7 million satoshis, commonly known as the “wumbo channel” threshold in the Lightning Network. Local indicates that the feature pertains to the outgoing edge, while remote refers to a feature of an incoming edge.
TABLE I: Features and their correlations to channel balance.
@@
Local Maximum HTLC 0.6246
Remote Maximum HTLC 0.6165
Remote Feature 19 is Known 0.1603
Local Feature 19 is Known 0.1523
Both Maximum HTLC and Feature 19 show a positive correlation with the target variable, as indicated by their positive correlation coefficients. Local Maximum HTLC and Remote Maximum HTLC exhibit a notably strong positive correlation. In contrast, Local Feature 19 is Known and Remote Feature 19 is Known exhibit a moderate yet meaningful positive correlation.
@@ VII-A Modeling
where $\Theta$ are learnable weights, $\mathbf{x}_{u}$, $\mathbf{x}_{v}$ and $\mathbf{e}_{(u,v)}$ are the node and edge features respectively, while $c_{\{u,v\}}$ is the channel capacity. While several choices are possible for $f_{\Theta}$, such as multi-layer perceptrons or Graph Neural Networks, we focus on Random Forests for this work given their simplicity and efficacy. In particular, our Random Forest (RF) model operates on the concatenation of the features of the source and destination nodes as well as the edge features:
[…]
The model is trained using a Mean Squared Error loss. Moreover, we know that $p_{(u,v)}=1-p_{(v,u)}$. Therefore, we would like to design our model such that $f_{\Theta}(u,v,G,x,c)=1-f_{\Theta}(v,u,G,x,c)$. We can obtain this approximately through data augmentation: make sure that both directions of an edge are present in the data so that the model is trained to predict both $f_{\Theta}(u,v,G,x,c)=p_{(u,v)}$ and $f_{\Theta}(v,u,G,x,c)=1-p_{(v,u)}$.
@@ VII-B Node Features
    The ratio of a node’s capacity to the total network capacity, indicating the proportion of network capacity connected to the node.
@@
    The ratio of a node’s average outgoing fee to its average incoming fee.
@@ VII-C Edge Features
  • Min HTLC The minimum amount this edge will route. (denominated in millisatoshis)
  • Max HTLC The maximum amount this edge will route. (denominated in millisatoshis)
  • Fee Rate Proportional fee to route a payment along an edge, measured in parts per million (ppm) satoshis. (1 ppm = 1 satoshi earned for every million satoshi routed).
  • Fee Base Fixed fee to route a payment along an edge. (denominated in millisatoshis)
@@ VII-D Positional Encodings
Positional encoding is essential for capturing the structural context of nodes, as graphs lack inherent sequential order. Using eigenvectors of the graph Laplacian matrix as positional encodings provides a robust solution to this challenge. These eigenvectors highlight key structural patterns about the overall topology of the graph [15]. By incorporating these spectral properties, machine learning models can better identify and leverage network-wide characteristics, improving performance in tasks such as node classification and community detection.
In practice, computing positional encodings for networks of the scale studied in this work ($7,497$ nodes and $40,462$ edges) takes approximately one second on a standard machine. While this computation could become more expensive on significantly larger graphs, the current performance is sufficient for near realtime use cases on today’s Lightning Network.